# CFBench: A Comprehensive Constraints-Following Benchmark for LLMs

### v1

@@ Introduction
CfBench 1000 10 [Uncaptioned image][Uncaptioned image]
@@
Data Quality
Quality Evolution
@@
Quality Evaluation
[…]
@@
Experiment
Evaluation Settings
We evaluated a total of 50 models that demonstrated exceptional performance on previous benchmarks  (Hendrycks et al. 2020; Cobbe et al. 2021), considering factors such as model size, support for Chinese among mainstream languages, whether they were accessed via API or weights, and whether they were fine-tuned with instruction tuning data. During inference, we set the maximum generation length to 2048, with other parameters using their default values. For evaluation, we used GPT-4o as the judge model, setting the temperature to 0 to ensure deterministic outputs. All model names mentioned correspond to the versions listed in Tab. 7 of the Appendix.
@@
Figure 6: Factors Influencing Constraints-Following Performance
Overall Result
Tab. 2 presents the detailed evaluation results of CFBench and its splits for the leading models, using three key metrics. Overall, GPT-4o leads across all splits and metrics, with Claude-3.5-Sonnet close behind. Qwen2-72B-Instruct, GLM4, DeepSeek-V2, and ERNIE4 also performed well, while Yi-large, abab6.5, and MoonShot lagged slightly. Considering the different splits, the Hard split shows a significant drop in metrics, highlighting the clear distinction between Hard and Easy splits. The highest PSR only reached 0.582, indicating room for improvement. Regarding the metrics, CSR is more favorable for weaker models, making it easier to identify issues at a fine-grained level. ISR and PSR highlight differences among stronger models, with ISR being the most challenging and PSR considering user priority and overall satisfaction. Notably, API-accessed models significantly outperformed open-source models accessed via weights, except for Qwen2-72B-Instruct, highlighting a substantial gap in constraint follow for open-source models.
Constraints-categorized Performance
To observe the performance across different constraint types, we calculated satisfaction scores for each constraint for the top 8 LLMs, as shown in Fig. 4. For primary constraints, many models struggled with contradictory constraints, revealing their limitations in handling conflicts. By considering requirement priorities, we captured nuances among multiple constraints, enabling detailed evaluation. GPT-4o excelled across various constraints, while other models alternated in leading different types. For secondary constraints, all models performed poorly in lexical, word, and sentence count constraints but did better in document count and audience style constraints. No single model consistently led across most constraint types. In summary, even the most advanced LLMs have significant room for improvement in certain constraint types, with each model showing specific weaknesses. This provides valuable insights for future model iterations.
Domain and Task-categorized Performance
@@
Discussion
Comparisons between Capabilities
Tab. 4 presents a comprehensive comparison of CFBench’s PSR with two of the prominent LLM evaluation benchmarks: MMLU (Hendrycks et al. 2020) and GSM8K (Cobbe et al. 2021). MMLU focuses on knowledge proficiency, while GSM8K emphasizes mathematical ability. GPT-4o ranks first on CFBench but significantly lags behind, ranking third on GSM8K. Qwen1.5-110B-Chat performs worse than DeepSeek-V2 on CFBench but outperforms it on MMLU. Notably, the rankings of LLMs on CFBench do not entirely correspond with those on the other two benchmarks, indicating that CFBench provides a novel perspective for LLMs evaluation.
Factors influencing constraints-following
@@ Discussion
How to improve constraint-following ability?
@@
In Appendix Tab. 7, we investigated methods to potentially enhance constraint following. Firstly, Supervised Fine-Tuning (SFT) significantly improves performance, with nearly all models that undergo instruction fine-tuning exhibiting substantial improvements in effectiveness, as demonstrated by the Qwen series. Secondly, model size is also an important factor, as evidenced by Qwen2-72B-Instruct showing a 40% relative PSR improvement over Qwen2-7B-Instruct. Additionally, replicating Conifer’s models  (Sun et al. 2024a) reveals that fine-tuning with complex constraint instructions further enhances performance, and recent work has also been directed towards this approach  (He et al. 2024a). Further exploration is intended to be pursued in future work.
Conclusion
@@
This study comprehensively examines the constraints-following capabilities of LLMs. CFBench, a comprehensive benchmark, was introduced with 1000 manually annotated samples covering more than 200 real-world scenarios and over 50 NLP tasks, encompassing a wide range of systematically defined constraint types. Each sample in CFBench includes detailed evaluation criteria, providing metrics that accurately reflect model performance and real user needs across multiple dimensions. Extensive experiments on CFBench revealed significant limitations and challenges that advanced LLMs face in following constraint instructions. Key factors and potential strategies to improve constraint following were also analyzed. In conclusion, CFBench offers a novel perspective for evaluating LLM capabilities, providing new directions for performance assessment and improvement.

### v2

@@ 3.4.2 Evaluation Metrics
FoFoBench 494 1 [Uncaptioned image]
ComplexBench 1150 4 [Uncaptioned image]
CFBench 1000 10-25 [Uncaptioned image][Uncaptioned image]
@@
3.5 Data Quality
3.5.1 Quality Evolution
@@
3.5.2 Quality Evaluation
To investigate CFBench’s quality, we randomly selected 100 samples for assessment. Three professional data inspectors evaluated them, resulting in high-quality rates of 94% for instructions, 94% for gold answers, and 93% for checklists (see Appendix Table 5). Additionally, three experts rated Qwen2-7B-Instruct outputs on a 0-1 scale. The kappa coefficient between GPT-4o PSR and expert evaluations was 0.77, highlighting the effectiveness of the PSR evaluation method and metrics, even for smaller models. Further details are in Appendix Table 6.
@@ 3.5.2 Quality Evaluation
4 Experiment
4.1 Evaluation Settings
We evaluated 50+ top-performing models from previous benchmarks Hendrycks et al. (2020); Cobbe et al. (2021), considering factors like model size, Chinese language support, access via API or weights, and fine-tuning with instruction data. During inference, we set the maximum generation length to 2048 and used default values for other parameters. For evaluation, we used GPT-4o as the judge model with a temperature of 0 for deterministic outputs.
Figure 6: Factors Influencing Constraints-Following Performance
4.2 Overall Result
Tab. 2 presents CFBench evaluation results for leading models. DeepSeek-R1 DeepSeek-AI et al. (2024) leads overall, followed by GPT-4o and DeepSeek-V3 DeepSeek-AI et al. (2025) in third. Claude-3.5-Sonnet Anthropic (2024) and Qwen2-72B-Instruct Yang et al. (2024) performed well, though DeepSeek-V3 and Qwen2-72B-Instruct showed slight drops in the Full Set. The highest PSR in the Hard Set was 0.582, indicating room for improvement.
While CSR favors weaker models, ISR and PSR highlight differences in stronger models. API-accessed models like GPT-4o outperformed most open-source models, though DeepSeek-V3 and Qwen2-72B-Instruct performed well among open-source models.
@@ 4.2 Overall Result
4.3 Constraints-categorized Performance
To assess performance across different constraint types, we calculated satisfaction scores for the top 8 LLMs (see Fig. 4). Many models struggled with contradictory constraints, highlighting their limitations. GPT-4o excelled across various constraints, while other models alternated in leading different types. For secondary constraints, all models performed poorly in lexical, word, and sentence count constraints but did better in document count and audience style constraints. No single model consistently led across most constraint types. In summary, even the most advanced LLMs have significant room for improvement, with each model showing specific weaknesses, providing valuable insights for future iterations.
4.4 Domain and Task-categorized Performance
@@
5 Discussions
5.1 Comparisons between Capabilities
Table 4 presents a comprehensive comparison of CFBench’s PSR with two prominent LLM evaluation benchmarks: MMLU Hendrycks et al. (2020) and GSM8K Cobbe et al. (2021). MMLU focuses on knowledge proficiency, while GSM8K emphasizes mathematical ability. GPT-4o ranks first on CFBench but significantly lags behind, ranking third on GSM8K. Qwen2-72B-Instruct performs worse than DeepSeek-V2 on CFBench but outperforms it on MMLU. Notably, the rankings of LLMs on CFBench do not entirely correspond with those on the other two benchmarks, indicating that CFBench provides a novel perspective for LLMs evaluation.
5.2 Factors influencing constraints-following
@@
5.3 How to improve constraint-following ability?
@@
In Appendix Tab. 7, we investigated methods to potentially enhance constraint following. Firstly, Supervised Fine-Tuning (SFT) significantly improves performance, with nearly all models that undergo instruction fine-tuning exhibiting substantial improvements in effectiveness, as demonstrated by the Qwen series. Secondly, model size is also an important factor, as evidenced by Qwen2-72B-Instruct showing a 40% relative PSR improvement over Qwen2-7B-Instruct. Additionally, replicating Conifer’s models  Sun et al. (2024a) reveals that fine-tuning with complex constraint instructions further enhances performance, and recent work has also been directed towards this approach  He et al. (2024a). Further exploration is intended to be pursued in future work.
6 Conclusion
@@
This study comprehensively examines the constraints-following capabilities of LLMs. CFBench, a comprehensive benchmark, was introduced with 1000 manually annotated samples covering more than 200 real-world scenarios and over 50 NLP tasks, encompassing a wide range of systematically defined constraint types. Each sample in CFBench includes detailed evaluation criteria, providing metrics that accurately reflect model performance and real user needs across multiple dimensions. Extensive experiments on CFBench revealed significant limitations and challenges that advanced LLMs face in following constraint instructions. Key factors and potential strategies to improve constraint following were also analyzed, and numerous insightful findings can provide valuable guidance for the optimization of LLMs’ performance. In conclusion, CFBench offers a novel perspective for evaluating LLM capabilities, providing new directions for performance assessment and improvement.
7 Limitations
7.1 Experimental Setup
This study primarily focuses on models with strong Chinese language capabilities, lacking a comprehensive survey of a broader range of English models. Additionally, while we conducted preliminary analyses on the differences in instruction-following abilities between Chinese and English, a more in-depth comparative study is absent.
7.2 Limited Exploration of Reasoning Models
Currently, deep reasoning models like R1 continue to achieve commendable results. However, there is a lack of in-depth research into these models, particularly concerning the factors that enhance their instruction-following abilities.
7.3 Evaluation Model Bias
The evaluation of models predominantly relies on GPT-4o as the judge model. Future research could explore the impact of different evaluation models on assessing the performance of other models.
8 Ethics Statement
This research adheres to the ethical guidelines set forth by the Association for Computational Linguistics (ACL). We have ensured that all data collection and experimental designs comply with privacy protection and informed consent principles, fully respecting and safeguarding the rights of all participants. Furthermore, we have evaluated the potential societal impacts of our research findings, ensuring that their application does not result in adverse effects on society.
Appendix A Appendix
A.1 Constraint System Construction
The construction of the constraint system commenced with the aggregation of data from diverse real-world scenarios and NLP tasks. This encompassed 800,000 query logs from LLM websites over the preceding six months, alongside over 300,000 data points from various NLP tasks. Instructions that were excessively long or short were filtered out, and a vector clustering deduplication algorithm was employed. This meticulous process culminated in a refined dataset comprising approximately 30,000 instructions. Subsequently, GPT was utilized to extract constraint atoms from these instructions, thereby ensuring the comprehensiveness of the constraint system. The prompt employed for GPT-4 extraction, as illustrated in Fig. 7, resulted in the identification of approximately 5,000 unique atomic constraint instructions. Three seasoned experts meticulously refined these into 1,000 significant atomic constraints. By integrating statistical analysis, classification, and linguistic principles, a hierarchical constraint system was developed using Top-Down Organization and Bottom-Up Synthesis methodologies. This system comprises 10 primary categories and 25 secondary categories. The system comprehensively categorizes all types of constraints, ensuring that nearly all specific constraints can be systematically classified within its framework. Detailed information regarding the constraint system is presented in Tab. 8.
A.2 Dataset Construction
We adopted an innovative Human-Machine Collaborative Iterative Construction approach to ensure the highest quality of data. This method involved leveraging advanced LLMs to augment original instructions with additional constraints and generate corresponding responses. These responses were meticulously reviewed for constraint validity, followed by the creation of detailed checklists for each example. Multiple experts participated in this iterative process, continuously refining the outputs by addressing issues encountered by the LLMs and regenerating or manually correcting any substandard samples. The prompts used for GPT to enhance constraints and generate checklists are illustrated in Fig. 8 and Fig. 10. Due to the limited attention given to real-life scenarios, we have meticulously organized and covered 20 domains and over 200 scenarios in our CFBench system, as detailed in Tab. 9. In the end, we gathered 1,000 high-quality data points: 500 from real-world scenarios and 500 from different NLP tasks. Specifically, we implemented the following steps to enhance data quality for manual annotations.
A.2.1 Annotator Training
We sourced annotation contractors from the public and selected 21 candidates for training by seasoned data scientists. After a one-week training program, the annotators engaged in multiple rounds of trial annotations, which were then assessed by data experts. From these assessments, the 9 annotators demonstrating the highest accuracy were selected for this dataset.
A.2.2 Cross-Validation
To reduce the likelihood of missed and incorrect annotations, we implemented an inter-annotator validation process. Three annotators independently reviewed the labeled instructions, responses, and evaluation criteria, achieving a notable agreement rate of 94%. Any discrepancies that emerged were resolved through expert adjudication, ensuring both consistency and accuracy.
A.2.3 Batch Validation
Due to the substantial size of the dataset, it was systematically divided for processing. Following a phased improvement approach, the initial batch sizes were set at 50, 100, and 200, gradually increasing to 400 for later batches. After the annotation process, 50% of the dataset was randomly selected for contractor review, while 20% of the dataset was examined by experts.
A.2.4 Data Split
We used a voting mechanism involving experts and ten models, including GPT-4o and Claude-3.5-Sonnet, to partition 10,000 CFBench entries into ’easy’ and ’hard’ categories. The ’hard’ category includes entries where multiple models struggle with PSR performance and are also challenging for humans, as verified by experts.
Figure 7: Prompt Template Atomic Constraint Extract
Figure 8: Prompt Template for Adding Atomic constraint
A.3 Evaluation Method and Metric
The state-of-the-art GPT-4o model was employed as the judge to perform binary scoring (0 or 1) for each checkpoint in the checklist. The specific evaluation prompt is illustrated in Fig. 9. The Requirement Priority-Satisfaction Ratio (PSR) was proposed as an evaluation metric that simultaneously considers the prioritization of user requirements and satisfaction levels. PSR is calculated by first ensuring that all primary requirements are met. Subsequently, the satisfaction score is determined by averaging the fulfillment of the remaining constraints to obtain $A$. The final satisfaction score is then calculated using the formula $0.5+0.5*A$. If the final score exceeds $0.8$, PSR is set to $1$. The threshold of $0.8$ was established based on the average satisfaction levels derived from multiple users’ feedback on the responses to the instructions.
A.4 Quality Assessment
We employed multiple methods to validate the quality of the benchmark on a randomly selected set of 100 samples. First, we engaged three experts to independently evaluate the quality of each sample’s instruction, response, and criteria. The average quality rate determined by the three experts was consistently above 90%, as detailed in Tab. 5. To further validate the effectiveness of our proposed evaluation metric, PSR, we had the same three experts score the responses of Qwen2-7B-Instruct on these 100 cases using a 0-1 scale. Simultaneously, we utilized GPT-4o to directly score the responses, referred to as GPT-4o PSR. By calculating the kappa coefficient, we found a strong agreement between our proposed PSR evaluation metric and the human experts’ assessments. The detailed results are presented in Tab. 6. Kappa coefficient scores are interpreted as follows: below 0.2 indicates slight agreement, 0.21 to 0.40 indicates fair agreement, 0.41 to 0.60 indicates moderate agreement, 0.61 to 0.80 indicates substantial agreement, and 0.81 to 1.00 indicates almost perfect agreement.
Figure 9: Prompt Template for Evaluation
Set Instruction Gold Ans CheckList
Easy Set 0.96 0.94 0.93
Hard Set 0.92 0.95 0.93
All Set 0.94 0.94 0.93
Table 5: The High-Quality Rate of 100 selected Samples
Set Easy Set Hard Set Full Set
Avg.Expert 1 1 1
GPT-4o DS 0.58 0.61 0.60
GPT-4o PSR 0.76 0.77 0.77
Qwen2-72B-Inst. PSR 0.70 0.73 0.72
Table 6: The kappa coefficient between expert evaluations and various assessment methods
A.5 Experimental Setup and Results
A.5.1 Experiment Setting
We evaluated the most popular Large Language Models (LLMs), with the majority of these models being developed by companies based in China, primarily to accommodate our CFBench’s focus on the Chinese language. Among the 50 evaluated models, they can be categorized into two groups based on their access method: API-based and open-source weight-based models. It is worth noting that the Llama series models do not primarily support the Chinese language, which results in noticeably lower performance. Both conifer-base and conifer-test are based on the Mistral-7B foundational model. Llama-3-8B-Instruct-CN and Llama-3-70B-Instruct-CN respectively represent Llama-3-8B-Instruct-Chinese and Llama-3-70B-Instruct-Chinese, both of which have undergone Chinese SFT (Supervised Fine-Tuning). For the base models, we used a 3-shot approach to ensure a fair evaluation. The complete list of evaluated models can be found in Tab. 7.
A.5.2 Explanation of Results
GPT-4o and Claude3.5-Sonnet have demonstrated near-absolute leadership, achieving outstanding performance across various metrics and categories. Similarly, models such as GLM-4-0510, ERNIE-4-Bot-0613, ERNIE-4-Turbo-0628, DeepSeek-V2-0628, and Qwen2-72B-Instruct have also exhibited strong capabilities. Many models that support less mainstream Chinese languages performed significantly worse, which is unfair to them and only serves to illustrate their relative rankings. This also confirms that performance are highly correlated with language, especially within the scope of language constraints. From the perspective of open-source versus closed-source models, open-source models have generally achieved comprehensive success. However, Qwen2-72B-Instruct, as an open-source model, also demonstrated notable constraint-following capabilities. Regarding model size, within the Qwen series, performance metrics clearly improve with increasing model size. Additionally, models that have undergone Supervised Fine-Tuning (SFT) show significantly enhanced instruction-following capabilities. The complete evaluation results and rankings can be found in Tab. 7.
Models Easy Set Hard Set Full Set
CSR ISR PSR CSR ISR PSR CSR ISR PSR
GPT-4o 0.956 0.868 0.888 0.816 0.438 0.582 0.886 0.653 0.735
GLM-4-0520 0.939 0.820 0.852 0.785 0.372 0.536 0.862 0.596 0.694
ERNIE-4-Bot-0613 0.929 0.792 0.832 0.779 0.338 0.518 0.854 0.565 0.675
ERNIE-3.5-0613 0.901 0.720 0.772 0.758 0.302 0.482 0.830 0.511 0.627
Yi-Large 0.900 0.730 0.786 0.744 0.292 0.460 0.822 0.511 0.623
abab6.5-chat 0.894 0.696 0.766 0.736 0.260 0.452 0.815 0.478 0.609
MoonShot-V1-8k 0.919 0.764 0.812 0.758 0.308 0.464 0.838 0.536 0.638
Vicuna-7B-V13 0.563 0.206 0.262 0.468 0.100 0.168 0.516 0.153 0.215
Vicuna-33B-V13 0.621 0.270 0.352 0.527 0.110 0.196 0.574 0.190 0.274
Vicuna-13B-V13 0.605 0.248 0.302 0.503 0.100 0.178 0.554 0.174 0.240
Llama-2-7B-Chat 0.5268 0.198 0.250 0.448 0.096 0.152 0.487 0.147 0.201
Llama-2-13B-Chat 0.574 0.242 0.280 0.488 0.094 0.178 0.531 0.168 0.229
Mistral-7B-Instruct-V03 0.227 0.072 0.086 0.148 0.008 0.022 0.188 0.040 0.054
Conifer-Base 0.510 0.184 0.232 0.300 0.018 0.048 0.405 0.101 0.140
Conifer-Test 0.559 0.215 0.255 0.328 0.102 0.156 0.443 0.159 0.206
BaiChuan-13B-Chat 0.630 0.306 0.366 0.521 0.114 0.196 0.575 0.210 0.281
BaiChuan2-13B-Chat 0.669 0.348 0.418 0.547 0.134 0.226 0.608 0.241 0.322
Llama-3-8B-Instruct-CN 0.743 0.458 0.510 0.627 0.162 0.314 0.685 0.310 0.412
Llama-3-70B-Instruct-CN 0.756 0.482 0.536 0.636 0.190 0.322 0.696 0.336 0.429
DeepSeek-7B-Chat 0.695 0.378 0.442 0.580 0.150 0.270 0.638 0.264 0.356
DeepSeek-67B-Chat 0.802 0.516 0.578 0.662 0.180 0.350 0.732 0.348 0.464
InternLM2-Chat-7B 0.767 0.452 0.538 0.625 0.172 0.320 0.696 0.312 0.429
GLM-4-9B-Chat 0.885 0.678 0.742 0.742 0.288 0.450 0.813 0.483 0.596
Qwen1.5-4B 0.454 0.170 0.198 0.376 0.074 0.116 0.415 0.122 0.157
Qwen1.5-4B-Chat 0.652 0.310 0.362 0.536 0.104 0.198 0.594 0.207 0.280
Qwen1.5-7B 0.473 0.176 0.212 0.400 0.090 0.142 0.437 0.133 0.177
Qwen1.5-7B-Chat 0.799 0.534 0.592 0.654 0.194 0.338 0.726 0.364 0.465
Qwen1.5-14B 0.498 0.228 0.280 0.430 0.110 0.176 0.464 0.169 0.228
Qwen1.5-14B-Chat 0.822 0.558 0.626 0.671 0.202 0.370 0.746 0.380 0.498
Qwen1.5-32B 0.647 0.336 0.408 0.528 0.132 0.224 0.587 0.234 0.316
Qwen1.5-32B-Chat 0.883 0.678 0.744 0.704 0.228 0.412 0.793 0.453 0.578
Qwen1.5-72B 0.627 0.324 0.380 0.556 0.148 0.248 0.591 0.236 0.314
Qwen1.5-72B-Chat 0.896 0.710 0.776 0.730 0.254 0.436 0.813 0.482 0.606
Qwen2-0.5B-Instruct 0.446 0.150 0.172 0.393 0.070 0.110 0.419 0.110 0.141
Qwen2-1.5B-Instruct 0.607 0.250 0.316 0.496 0.104 0.168 0.551 0.177 0.242
Qwen2-7B 0.576 0.260 0.316 0.478 0.120 0.192 0.527 0.190 0.254
Qwen2-7B-Instruct 0.835 0.584 0.642 0.682 0.198 0.362 0.758 0.391 0.502
Qwen2-72B 0.711 0.424 0.484 0.568 0.170 0.274 0.640 0.297 0.379
Table 7: The complete evaluation results and rankings of CFBench and its respective subsets. Notably, stands for the model supporting mainstream languages excluding Chinese, and represents calling through the API. The bold, underlined, and tilde denote the first, second, and third rankings, respectively. Llama-3-8B-Instruct-CN and Llama-3-70B-Instruct-CN respectively represent Llama-3-8B-Instruct-Chinese and Llama-3-70B-Instruct-Chinese, both of which have undergone Chinese SFT (Supervised Fine-Tuning). Both coniferbase and conifertest are based on the Mistral-7B foundational model. For the base model, we used a 3-shot approach for generation.
Primary Secondary Definition Example
Content Constraint Lexical Mandatory use of specific terms or symbols, including their inclusion and precise placement. …must include the word "beautiful."
Element Mandates for including specific elements or concepts in responses, reflecting a scenario or object. …highlights the Great Wall.
Semantic Directives on thematic content, perspective, or tone, emphasizing response significance. Write a poem about London.
Numerical Constraint Word Count Limit the number of words or tokens. A 50-word poem.
Sentence Count Limit the number of sentences. … three sentences.
Paragraph Count Limit the number of paragraphs. divided into 3 sections.
Document Count Limit the number of documents. … list 3 articles.
Stylistic Constraint Tone and emotion The emotional tone must adhere to standards such as seriousness, anger, joy, humor, and politeness. Write a letter in an angry and sarcastic tone.
Form and style Text expression standards ensure alignment with specific stylistic criteria in both presentation and perception. Write a passage in an encyclopedic style.
Audience-specific Text should be tailored to specific audiences, ensuring clarity and relevance for children, students, or specialized groups. Write a pome for a 6-year-old.
Authorial style Texts should emulate the styles of authors like Shakespeare to achieve artistic effects or depth. Write a passage the style of Shakespeare.
Format Constraint Fundamental Widely accepted and utilized standard formats, including JSON, XML, LaTeX, HTML, Table, and Markdown. Extract keywords and output in JSON format.
Bespoke Protocols for information expression tailored to specific needs,including paragraphing, headings, text emphasis, examples, and bullet points. Summarize the main idea and output in unordered list format.
Specialized Formatting standards tailored for specialized applications or domains. Conform to electronic medical record format.
Linguistic Constraint Pragmatic Contextual language study, encompassing speech acts, implicature, discourse, dialects, sociolects, and language policy. Output in English, in classical Chinese style.
Syntactic Sentence structure, including phrases, constituents, subordinate clauses, ba-constructions, and imperatives. Use imperatives with nouns and verb phrases.
Morphological The internal structure and formation rules of words, including roots, affixes, and morphological changes. Output all content in lowercase English.
Phonological Study on phonological structures:phonemes, allophones, pitch, duration, and intensity. Single-rhyme tongue twisters.
Situation Constraint Role-based Simulating characters based on context, emulating their traits, language, and behaviors. You are Confucius, how do you decide?
Task-specific Offer tailored solutions based on a nuanced understanding of situational demands. Must work from home, how to report?
Complex context Reasoning and problem-solving within intricate and multifaceted contexts. 4 on the left, 10 total, which from right?
Example Constraint - Regulate new responses by leveraging intrinsic patterns from a limited set of samples. Example:input:xxx, output:{…}; input:xx, output?
Inverse Constraint - Narrow the response space through inverse constraints and indirect exclusion. Prohibited from answer political topics.
Contradictory Constraint - Mutually exclusive constraints prevent fulfilling all requirements concurrently. Write a five-character quatrain, 1000 words.
Rule Constraint - Standardize the road of responses through meticulously crafted logic flows or actions. Each answer adds 1, 1+1=3, then 2+3=?
Table 8: Constraint System of CFBench
Domain Scenarios List
Healthcare Symptom Consultation Diagnostic Explanation Medication Guidance Procedures
Wellness Medical Info Guidelines Inquiry Public Health
Medical Education
Education Teaching Methods Resource Access Curriculum Design Communication
Academic Counseling Mental Health Support Tutoring Subject Q&A
Reports Interests
Finance Market Research Stock Analysis Investment Analysis Personal Finance
Corporate Tax Insurance Management Corporate Financing Compliance & Risk
Financial Education Product Development Customer Service Financial Reports
Regulatory Analysis
Legal Legal Education Legal Consultation Document Review Case Analysis
Statute Explanation Regulation Analysis IP Management Legal Training
Case Management Compliance & Risk
Media Content Creation Information Analysis Marketing & Promotion News Reporting
Tourism Travel Consultation Itinerary Planning Route Introduction
Recruitment JD Writing & Analysis Resume Creation Resume Screening Interview Preparation
Interview Evaluation Career Planning Offer Comparison Communication Skills
Performance Review
Gov Affairs Policy Research Public Education Service Guide Public Services
Document Writing Content Review Business Procedures Civil Servant Training
Emergency Management
Real Estate Purchase Planning Market Trends Property Policies Development
Leasing Property Valuation Amenities Financial Services
Property Description Content Creation Sales & Marketing Qualifications
Renovation
Automotive Marketing & Sales Driving & Safety Customer Experience Model Consultation
Model Comparison Loan Calculation Insurance Evaluation Claims Assessment
Car Reviews Maintenance & Repair
Psychology Romance Family Friendship Workplace
Self & Health Social Sexuality & Gender Life Stages
Organizational Client Relations Crisis Intervention Public Psychology
Internet Business Analysis Product Design User Research Coding & Debugging
Product Testing Data Management Cybersecurity Computer Q&A
Internet News Marketing Operations UI/UX Design
Spirituality Beliefs & Rituals Divination Feng Shui Astrology
Metaphysics Spirituality Healing Content Review
Sports Training Goal Setting Nutrition Workout Plans
Equipment & Tech Performance Injury Care Specialized Training
Mental Motivation Data Tracking
Lifestyle Life Tips Shopping Decisions Instant Queries Skincare
Fashion & Styling Naming Recommendations Planning
Socializing Life Creations Q&A
Culturtainment Podcasts & Radio TV & Film Music Literature
Theater & Dance Art Cultural Events Short Videos & Live
Gossip Content Creation
Employment Project Management Translation Office Efficiency Marketing
Administration Customer Service Team Collaboration Collaboration
Cuisine Food & Restaurant Recs Reviews & Feedback Marketing & Promo Food Content
Culinary Culture Recipes & Menus Cooking Techniques Ingredient Prep
Nutrition & Health Food Safety Culinary Training
Gaming Guides Reviews Hardware & Peripherals News
Software & Services Development & Design Operations Mini Games
Search Marketing & Promotion Esports & Tournaments Culture & Education
Technology Reviews Launches Buying Guides Tips & Tricks
Content Creation Product Design Marketing Copy After-sales & Repairs
Table 9: Domain and Scenarios List
Benchmark Case Features CFBench’s Advantages
Complex
Bench
Instruction: Please introduce the following painting.
- If the work contains any animal, the description should be in English.
- Otherwise, the description should be in Chinese.
Response: Painting: "Mona Lisa"
Focus on the combination of instructions, including And/Chain/Selection/Nested.
CFBench supports simplicity in instruction compliance, broad applicability across use cases, and scalability for various tasks.
Follow
Bench
Instruction: Using following words to generate a story.
"needed, money, computer, bought, happy".
Response: In the heart of the city, there was a research facility that desperately needed funding for a crucial project. To gather the required money, ···
Focus on gradual constraint escalation. The more constraints per instruction, the greater the complexity.
CFBench is simple to use, resource-efficient, and scalable for large models.
IFEval Instruction: Given the sentence "Two young boys with toy guns and horns." can you ask a question? Please ensure that your response is in English, and in all lowercase letters. No capital letters are allowed.
Response: what are the two young boys holding?
Verifiable Instructions.
CFBench is easy to implement, general-purpose across models, and scalable for large datasets.
CFBench Instruction: Overview of the Trump Shooting Incident, include the cause, process, aftermath, shooter information and Trump’s feedback. Organize the content using numbered points and bold text. The cause should be in JSON format with event keywords: time, location, and nature. Summarize the process in three sentences. Discuss the impact from both international and domestic perspectives, keeping the total content under 500 words. Finally, summarize this event in a seven-word quatrain.
[…]
Comprehensive Coverage:
Multi-dimensional constraints;
Cross-task evaluation;
Real-world complexity.
Standardized Framework:
Unified metrics;
Cross-model compatibility.
[Uncaptioned image] [Uncaptioned image]
Diverse instruction types,
standardized evaluation,
hierarchical constraints.
Table 10: Benchmark Comparison for Different Instruction Types and Task Complexity
Figure 10: Prompt Template for CheckList Generation
Figure 11: CFBench Example: English-Chinese Comparison, Data Itself in Chinese.